The perimeter of a rectangle is 36 and the length is twice the width. What are the dimensions of this rectangle?
Answer by macston(5194)

W=width; L=length=2W
.
P=2(L+W)
36=2(L+W)
18=L+W Substitute for L.
18=2W+W
18=3W
6=W ANSWER 1: Width is 6 units.
L=2W+2(6)=12 ANSWER 2: Length is 12 units
.
CHECK:
P=2(L+W)
36=2(12+6)
36=2(18)
36=36